OpenStudy (anonymous):

Need help w/ indefinite integral: Integral (3/x^5) from 0 to 1

OpenStudy (anonymous):

\[\int\limits_{0}^{1} (3/x^5) dx \]

OpenStudy (zarkon):

integral does not converge

OpenStudy (anonymous):

so for my final answer I ended up with: 3ln(1)-3ln(0)=0, not sure if that is correct...

OpenStudy (zarkon):

ln is not defined at 0

OpenStudy (anonymous):

so would it be 3ln(1)-3ln(t)^5 =0 ?

OpenStudy (zarkon):

no

OpenStudy (zarkon):

\[\int3/x^5dx=\frac{-3}{4x^4}+c\]

OpenStudy (zarkon):

there are no ln's in this problem

OpenStudy (anonymous):

ah, okay I see. Thanks